please solve and explain detailed
sinx + cosx= 1

sin (x) + cos(x) = 1
Answer #1

solution;- Given Güven trigonometric ege trigonometric equation is Bina +COS(O)=) Sina = 1 - COST Squaring on both sides (sincase 2 |- Casa = 0 -1082=-) COS Y = ) Cosa = cos (0) if cos O = cos as then o - 2012 az an TŁO - ANT solution of the given tr